Five bottles of milk cost $4.4 more than four bottles of juice. If one bottle of milk costs $5.2 find the cost of one bottle of juice
Answer by srinivas.g(540)

cost of one bottle of milk = $5.2
cost of five bottles of milk = 5*5.2
= $26
Five bottles of milk cost $4.4 more than four bottles of juice
cost of 4 bottle of juice =26-4.4
= $21.6
cost of 1 bottle of juice = 21.6/4
= $5.4